Suppose that King Kanuta, whom you met in your workbook, demands that each of his subjects give him 2 coconuts for every coconut that they consume. The king puts all of the coconuts that he collects in a large pile and burns them. The supply of coconuts is given by S(p.) 100ps, where p, is the price received by suppliers. The demand for coconuts by the kings subjects is given by Díp) -6,000 100p4 where pa is the price paid by consumers. In equilibrium, the price received by suppliers will be
Answer #1

At equilibrium,

quantity demanded = quantity supplied

S(PS) = D(Pd)

100*p = 6000-100*p

200*p = 6000

So p = 6000/200 = 30

So at equilibrium price is 30